Calcium carbonate decomposes at high temperatures to form carbon dioxide and calcium oxide. What mass of CaCO3 would be needed to produce 3.45L of CO2 at 1.0 atm and 289K?


1
Expert's answer
2021-10-28T02:57:35-0400

CaCO3 → CO2 + CaO

"V=3.45 \;L \\

p= 1.0 \;atm \\

T = 289 \;K"

Ideal gas law

pV =nRT

R = 0.08206 L×atm/mol×K

"n = \frac{pV}{RT} \\

n(CO_2) = \frac{1.0 \times 3.45}{0.08206 \times 289} = 0.145 \;mol"

According to the reaction:

n(CaCO3) =n(CO2) = 0.145 mol

MM(CaCO3) = 100.08 g/mol

m(CaCO3) "= 0.145 \times 100.08 = 14.56 \;g"

Answer: 14.56 g